5 Report your 1H NMR analyses here: (10 points) 6. Conclusion: (5 points) Draw Structure showing all carbons and hydrogens ani need help with this HNMR practice problem. i obtained an IHD of 6 but im unsure what this means as well.

2964 1729 43 2 68 O 6E 7 113 52 37 42 4 2. DU (IHD): Show calculation(2 pts) Comments on IHD (1pt) 3. Report IR analysis here. Tabulate (5 pts) Region Functional group Base Value Absorption, cm-1 (Exact number read from IR Give as a range. Read from correlation chart ectra 4000-3000 cm 3000-2500 2500-1700 cm 1700-1500 cm 1500-400 cm 4 Explain any IR special features in this box (2 pts) (Symmetry, conjugation, resonance, H-bonding, Evidence for Ortho, para, meta disubstitution or mono substitution.), if applicable
5 Report your 1H NMR analyses here: (10 points) 6. Conclusion: (5 points) Draw Structure showing all carbons and hydrogens and match the peaks with the protons. 7. Explain any special features on the back of this HNMR: symmetry, Homotopic, heterotopic, constants, etc. (2 points) CH untroun 37 80 7.70 7.60 7.50 3.0 ppm 01)
Answer #1

Dear candidate,

please go through attached image.

Q2)

DU=CnH2n+2

Lets consider our MF: C10H10O4

Now apply this value of 'n' in our formula,

10X2+2=22

our compound have 10 hydrogen atoms so,

(22-10)/2

=12/2

=6

Degrees of unsaturation is 6.

Q3)

IHD=(2n+2- no of hydrogen in given molecule)/2

=(2X10+2-10)/2

=(22-10)/2

=12/2

=6

Region Absorption cm-1 Functional group Fragments
4000-3000cm-1 3029 C-H stretch alkene C-H stretch
3000-2500cm-1 2964 C-H stretch alkane C-H stretch
2500-1700cm-1 1729 Ester C=O
1700-1500cm-1 1600 C=C C=C
1500-400cm-1

  

Sr No Chemical Shift Integration Splitting No of neighbours Fragments
1 3.82 3H Singlet 0 CH3-O-
2 7.56 2H dd 1 Aromatic
3 7.66 2H dd 1 Aromatic

7.66 3.82 7.56 7.56 3.82 7.66